# Refined Risk Bounds for Unbounded Losses via Transductive Priors

### v1

@@ 6.2 Polynomial time algorithm for batch logistic regression
$=\mathbb{P}\left(\langle X,{\theta^{\star}}\rangle\langle X,\theta\rangle<0\right)-2\mathbb{E}\left[\sigma(-\left\lvert\langle X,\theta\rangle\right\rvert)\mathbbm{1}\left(\langle X,{\theta^{\star}}\rangle\langle X,\theta\rangle<0\right)\right]+\mathbb{E}\left[\sigma(-\left\lvert\langle X,\theta\rangle\right\rvert)\right]$
@@
$\mathbb{P}\left(\langle X,{\theta^{\star}}\rangle\langle X,\theta\rangle<0\right)$ $\leq\mathrm{err}_{{\theta^{\star}}}(\theta)-\mathbb{E}\left[\sigma(-\left\lvert\langle X,\theta\rangle\right\rvert)\right]+2\mathbb{E}\left[\sigma(-\left\lvert\langle X,\theta\rangle\right\rvert)\mathbbm{1}\left(\langle X,{\theta^{\star}}\rangle\langle X,\theta\rangle<0\right)\right]$
$\leq\mathrm{err}_{{\theta^{\star}}}(\theta)+\mathrm{err}_{{\theta^{\star}}}({\theta^{\star}}).$
Now we note that by Assumption 6.2,
$\mathrm{err}_{{\theta^{\star}}}({\theta^{\star}})$ $=\mathbb{E}\left[\sigma(-\left\lvert\langle X,{\theta^{\star}}\rangle\right\rvert)\right]$
$\leq\mathbb{P}\left(\left\lvert\langle X,{\theta^{\star}}\rangle\right\rvert\leq T\right)+\sigma(-T)$
$\leq\mathbb{P}\left(\left\lvert\langle x,{\theta^{\star}}/\left\|{\theta^{\star}}\right\|\rangle\right\rvert\leq 1/T\right)+\frac{1}{1+e^{T}}$
$\leq\frac{d}{T}.$
@@
By the risk bound for $\mathrm{SVM}$ (Bousquet et al., 2020, Theorem 12), we have, with probability at least $1-d/T$, that
$\widetilde{\mathrm{err}}_{{\theta^{\star}}}(\hat{\theta})\leq\frac{C^{\prime}(d+\log T)}{T}$
@@
$\mathrm{err}_{{\theta^{\star}}}(\hat{\theta})$ $=\mathbb{E}\left[\mathbbm{1}(\langle X,\theta\rangle<0)\right]$
$\leq 2\mathbb{E}\left[\mathbbm{1}(\langle X,\theta\rangle<0)\mid{}Y=\operatorname{sign}(\langle X,{\theta^{\star}}\rangle)\right]$
$=2\widetilde{\mathrm{err}}_{{\theta^{\star}}}(\hat{\theta}).$
This concludes the proof. ∎
Finally, whenever $\|{\theta^{\star}}\|$ is large enough, and the $\hat{\theta}$ is close enough to ${\theta^{\star}}$, the clipped linear predictor induced by $\hat{\theta}$ admits a good performance as shown by the following Lemma 6.3.
Lemma 6.3.
@@
Proof of Lemma 6.3 We have
$\mathbb{E}_{(X,Y)\sim\mathcal{D}_{{\theta^{\star}}}}\left[-\log k_{\hat{\theta}}(X,Y)\right]$ $\leq\mathbb{E}_{(X,Y)\sim\mathcal{D}_{{\theta^{\star}}}}\left[-\log k_{\hat{\theta}}(X,Y)\mathbbm{1}\left(\lvert\langle X,{\theta^{\star}}\rangle\rvert\leq T\right)\right]$
@@
where the first inequality is by division of events and the second inequality is by $-\log k_{\hat{\theta}}(X,Y)\leq\log(2T)$. We first estimate the probability in the slab by Assumption 6.2,
@@
Proof of Theorem 6.1 Denote $(X_{T+1},X_{T+1})=(X,Y)$. We consider two cases. For the first case, suppose $\left\|{\theta^{\star}}\right\|\leq T^{4}$. We have by the exponential weights
@@
Furthermore, we have by Proposition 3.1
$-\sum\limits_{i=1}^{T+1}\log\left(\mathrm{EWA}_{K,\alpha}(X_{i},Y_{i})\right)$ $\leq-\sum\limits_{t=1}^{T+1}\log(\sigma(Y_{t}\langle X_{t},\theta^{\star}\rangle))+\alpha({\theta^{\star}})^{\top}\left(\sum\limits_{t=1}^{T+1}X_{t}X_{t}^{\top}\right){\theta^{\star}}+\frac{d}{2}\log\left(1+\frac{1}{8\alpha}\right)$
@@
$\leq\mathbb{E}_{(X,Y)}[-\log\sigma(Y\langle X,{\theta^{\star}}\rangle)]+\alpha\lambda_{\max}\left(\mathbb{E}[XX^{\top}]\right)\|{\theta^{\star}}\|^{2}+\frac{3d\log T}{T+1}$
$\leq\mathbb{E}_{(X,Y)}[-\log\sigma(Y\langle X,{\theta^{\star}}\rangle)]+\frac{4d\log T}{T},$
where the last inequality uses Assumption 6.1. For the second case, suppose $\|{\theta^{\star}}\|>T^{4}$. For this, we have
@@
$\leq\mathbb{E}\left[-\log\left(k_{\hat{\theta}}(X,Y)\right)\left(\mathbbm{1}\left(\exists t,Y_{t}\neq\operatorname{sign}(\langle X_{t},{\theta^{\star}}\rangle)\right)\right.\right.$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbbm{1}\left(\forall t,Y_{t}=\operatorname{sign}(\langle X_{t},{\theta^{\star}}\rangle),\mathbb{P}_{X,Y\sim\mathcal{D}_{{\theta^{\star}}}}\left(\langle X,{\theta^{\star}}\rangle\langle X,\hat{\theta}\rangle<0\right)>\frac{C_{1}(d+\log T)}{T}\right)$
$\left.\left.\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbbm{1}\left(\mathbb{P}_{X,Y\sim\mathcal{D}_{{\theta^{\star}}}}\left(\langle X,{\theta^{\star}}\rangle\langle X,\hat{\theta}\rangle<0\right)\leq\frac{C_{1}(d+\log T)}{T}\right)\right)\right]$
@@
Then, by Lemmas 6.3, 6.1 and 6.2, we have
@@ 7 Directions of Future Research
  • Our online-to-batch conversion argument in Section 6 only provides in-expectation results in important applications, such as logistic regression. While recent martingale-based arguments allow for high-probability excess risk bounds (see, e.g., (van der Hoeven, Zhivotovskiy, and Cesa-Bianchi, 2023)), these existing arguments do not extend straightforwardly to the transductive setup, where the natural martingale structure is not present. An even more challenging question would be to provide computationally efficient methods for logistic regression with no assumptions on the design, for which the excess risk bound of Corollary 6.1 holds with high probability.
@@
Acknowledgements
JQ and AR acknowledge support from the Simons Foundation and the NSF through award DMS-2031883.
@@ Appendix A Additional Results for Sparse Regression
Before discussing these discrete priors, we note that despite the minor penalty of the logarithm of the inverse smallest scaled singular value constant $\kappa_{s}$, there is a significant advantage: the predictor in Theorem 2.2 has an explicit Gaussian-type integral form. On the other hand, the predictor derived from sparsity priors can be quite challenging to exploit, as it might require too many samples from this prior (or equivalently, sampling from the posterior distribution $\rho_{t}$ is hard). However, from a practical implementation standpoint, it can still be effective, as claimed by simulations in (Rigollet and Tsybakov, 2011, 2012).
@@
where $H_{d}=\sum\nolimits_{i=1}^{d}\exp(-i)$. Although this prior is not data-dependent, it can be naturally combined with the data-dependent prior (6) by introducing a product prior that first samples the sparsity pattern vector according to (43), then uses (6) restricted to the coordinates corresponding to this sparsity pattern. Finally, the exponential weights algorithm is run with respect to this data-dependent prior. To slightly simplify the analysis, we will use a version of Theorem 2.1, where the additional clipping step is performed, and then we use aggregation using the exp-concavity of the squared loss in the bounded case.
@@
$\widetilde{\theta}_{t,p}=\arg\min_{\theta\in\Theta_{p}}\left\{\sum_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}+\lambda\sum\limits_{i=1}^{T}(\langle x_{t},\theta\rangle)^{2}\right\},$
@@
The clear limitation of the discrete prior approach is that sampling from (44) is challenging compared to evaluating an explicit predictor (8). The work of Rigollet and Tsybakov (2012) suggests a Metropolis–Hastings type algorithm for such sampling, though in the worst case its convergence rate can be quite slow.
@@
Proof. Fix any $p\in\mathcal{P}$ such that $|p|=s$. When restricted to the coordinates corresponding to the set $p$, the bound of Corollary C.3 in Appendix C gives the regret bound
@@
It remains to notice that for any $p\in\mathcal{P}$ such that $|p|=s$, the prediction $\operatorname{clip_{m}}\left(\langle x_{t},\widetilde{\theta}_{t,p}\rangle\right)$ is absolutely bounded by $m$. Finally, combining the exp-concavity of the squared loss in the bounded case (Vovk, 2001, Remark 3) (thus, we choose $\eta=\frac{1}{8m^{2}}$ in the exponential weights algorithm) and the fact that the predictor (44) is a posterior average with respect to the exponential weights algorithm with prior $\mu$ over $\mathcal{P}$, and by Lemma 2.2 with $\log\left(\frac{1}{\mu(p)}\right)=2s\log\left(\frac{ed}{s}\right)+\frac{1}{2}$ by (Rigollet and Tsybakov, 2012, Inequality (5.4)), we obtain that for any $s$-sparse $\theta^{\star}\in\mathbb{R}^{d}$,
@@ Appendix B Mixability of the Squared Loss for Unbounded Sets of Predictors
Lemma B.1 (Restatement of Lemma 2.3).
@@
satisfies $\frac{\partial^{2}h(y)}{\partial y^{2}}\geq 0$, making $h$ convex. Importantly, the derivation of $\frac{\partial^{2}h(y)}{\partial y^{2}}$ does not impose any boundedness condition on $f_{\theta}(x)$. Since $h$ is convex, its maximum is attained at either $-1$ or $1$. Therefore, by symmetry, $\widehat{f}$ satisfies
@@ Appendix C Revisiting Existing Bounds for Sequential Linear Regression
Our loss function is the logarithmic function, and we simply compute the total error for the mix-loss as suggested by Lemma 2.1. This straightforward computation is sufficient to recover all existing sequential online bounds, including the one in (Vovk, 2001), the bound for clipped ridge regression in (Kivinen and Warmuth, 1999), and the bound of Gaillard et al. (2019), which has already been discussed in Section 2. In the following, we discuss how essentially the same strategy, specific to sequential ridge regression, was previously employed by Zhdanov and Vovk (2009). Here, we mainly provide a detailed exposition of this approach and extend it to our transductive setup.
@@
Proposition C.1 (Generalization of (Azoury and Warmuth, 2001; Zhdanov and Vovk, 2009)).
@@
Although the bound of Proposition C.1 can be derived directly using linear algebraic arguments (which is essentially done in the proof of Theorem 4.6 in (Azoury and Warmuth, 2001)), we believe that the following proof strategy is more insightful. It provides this identity without prior knowledge of its possible existence. We note that the proof strategy and explicit identity in Proposition C.1 closely align with previous work, specifically with Zhdanov and Vovk (2009, Theorem 1), which establishes this identity for $A=I_{d}$ and acknowledges that a similar identity appears in Azoury and Warmuth (2001). Additionally, the work by Zhdanov and Kalnishkan (2013) explores sequential formulations of kernel ridge regression with the focus on similar identities. Although this formulation is not novel, it is included here for completeness and to support the transductive prior framework used in this paper. The extension to a general positive definite matrix $A$ as the regularizer corresponds directly to the transductive priors considered, simplifying derivations of bounds such as that in Gaillard, Gerchinovitz, Huard, and Stoltz (2019), as well as additional technical results used throughout the paper.
@@
$-\log\left(\mathbb{E}\exp(-(y-Z)^{2}))\right)=\frac{(\nu-y)^{2}}{2\sigma^{2}+1}+\frac{1}{2}\log(2\sigma^{2}+1).$
Proof. Using Lemma 3.3, we have
[…]
@@
$=-\log\left(\frac{1}{\sqrt{2\sigma^{2}+1}}\exp\left(-\frac{(\nu-y)^{2}}{2\eta\sigma^{2}+1}\right)\right)$
@@
We are ready to provide a proof of Proposition C.1.
@@
We want to compute the right-hand side of (11) for the class (46). The first computation uses Lemma 3.3. We have
@@
Another perspective on (47) is the following. We assume that at step $t$ we predict with the distribution $\rho_{t}$ over $\mathbb{R}^{d}$ using the exponential weights over the class (46) with respect to the logarithmic loss. By Lemma 3.3, we have
$-\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\limits_{t=1}^{T}\pi(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right)=-\sum\limits_{t=1}^{T}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\exp\left(-\pi(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right).$ (49)
@@
Clearly, the distribution $\rho_{t}$ is multivariate Gaussian. Using Lemma 3.3, we can immediately obtain
@@
Observe that if $\theta\sim\rho_{t}$, then $\langle x_{t},\theta\rangle\sim\mathcal{N}\left(\langle\widehat{\theta}_{t},x_{t}\rangle,\frac{1}{2\pi}x_{t}^{\top}\left(\sum\limits_{i=1}^{t-1}x_{i}x_{i}^{\top}+A\right)^{-1}x_{t}\right)$, which implies by Lemma C.1 that the following holds
@@
Finally, we have by Lemma C.2
@@
$-\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\limits_{t=1}^{T}\pi(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right)=\sum\limits_{t=1}^{T}\left(\pi(1-h_{x_{t}})(\langle\widehat{\theta}_{i},x_{t}\rangle-y_{t})^{2}+\frac{1}{2}\log\left(\frac{1}{1-h_{x_{t}}}\right)\right).$ (50)
Denote $A_{t}=x_{t}^{\top}\left(\sum\limits_{i=1}^{t-1}x_{i}x_{i}^{\top}+A\right)^{-1}x_{t}$. Finally, by Lemma C.2 we have
$\sum\limits_{t=1}^{T}\frac{1}{2}\log\left(\frac{1}{1-h_{x_{t}}}\right)=\frac{1}{2}\sum\limits_{t=1}^{T}\log\left(\frac{\det(A_{t+1})}{\det(A_{t})}\right)=-\frac{1}{2}\log(\det(A))+\frac{1}{2}\log\left(\det\left(\sum\limits_{t=1}^{T}x_{t}x_{t}^{\top}+A\right)\right).$ (51)
where in the last computation, we used the telescopic sum. Combining (48), (50), and (51), we prove the claim. ∎
The result of Proposition C.1 immediately implies the following classical result.
Corollary C.1 (The regret of the non-linear predictor of Vovk (2001)).
@@
Proof. To see how the implication, one should first verify, using Lemma C.1, that
@@
Combining this with Proposition C.1, we have

### v3

@@ 6.2 Polynomial time algorithm for batch logistic regression
[…]
@@
$\begin{split}&\mathbb{P}\left(\langle X,{\theta^{\star}}\rangle\langle X,\theta\rangle<0\right)\\ &\leq\mathrm{err}_{{\theta^{\star}}}(\theta)-\mathbb{E}\left[\sigma(-\left\lvert\langle X,\theta\rangle\right\rvert)\right]+2\mathbb{E}\left[\sigma(-\left\lvert\langle X,\theta\rangle\right\rvert)\mathbbm{1}\left(\langle X,{\theta^{\star}}\rangle\langle X,\theta\rangle<0\right)\right]\\ &\leq\mathrm{err}_{{\theta^{\star}}}(\theta)+\mathrm{err}_{{\theta^{\star}}}({\theta^{\star}}).\end{split}$ (41)
Now we note that by 30,
$\begin{split}\mathrm{err}_{{\theta^{\star}}}({\theta^{\star}})&=\mathbb{E}\left[\sigma(-\left\lvert\langle X,{\theta^{\star}}\rangle\right\rvert)\right]\\ &\leq\mathbb{P}\left(\left\lvert\langle X,{\theta^{\star}}\rangle\right\rvert\leq T\right)+\sigma(-T)\\ &\leq\mathbb{P}\left(\left\lvert\langle x,{\theta^{\star}}/\left\|{\theta^{\star}}\right\|\rangle\right\rvert\leq 1/T\right)+\frac{1}{1+e^{T}}\\ &\leq\frac{d}{T}.\end{split}$ (42)
@@
By the risk bound for $\mathrm{SVM}$ (Bousquet et al., 2020, Theorem 12), we have, with probability at least 1d/T1-d/T, that
$\widetilde{\mathrm{err}}_{{\theta^{\star}}}(\hat{\theta})\leq\frac{C^{\prime}(d+\log T)}{T}$ (43)
@@
[…]
Combining Equations 41, 42, 43 and 44 concludes the proof. ∎
Finally, whenever $\|{\theta^{\star}}\|$ is large enough, and the $\hat{\theta}$ is close enough to $\theta^{\star}$}, the clipped linear predictor induced by $\hat{\theta}$ admits a good performance as shown by the following Lemma 33.
Lemma 33.
@@
Proof of Lemma 33 We have
$\hskip-72.26999pt\mathbb{E}_{(X,Y)\sim\mathcal{D}_{{\theta^{\star}}}}\left[-\log k_{\hat{\theta}}(X,Y)\right]$
$\leq\mathbb{E}_{(X,Y)\sim\mathcal{D}_{{\theta^{\star}}}}\left[-\log k_{\hat{\theta}}(X,Y)\mathbbm{1}\left(\lvert\langle X,{\theta^{\star}}\rangle\rvert\leq T\right)\right]$
@@
where the second inequality is by $\log k_{\hat{\theta}}(X,Y)\leq\log(2T)$. We first estimate the probability in the slab by 30,
@@
Proof of Theorem 28 Denote $(X_{T+1},X_{T+1})=(X,Y)$. We consider two cases. For the first case, suppose $\left\|{\theta^{\star}}\right\|\leq T^{4}$. We have by the exponential weights
@@
Furthermore, we have by Proposition 15
$\hskip-72.26999pt-\sum\limits_{i=1}^{T+1}\log\left(\mathrm{EWA}_{K,\alpha}(X_{i},Y_{i})\right)$
$\leq-\sum\limits_{t=1}^{T+1}\log(\sigma(Y_{t}\langle X_{t},\theta^{\star}\rangle))+\alpha({\theta^{\star}})^{\top}\left(\sum\limits_{t=1}^{T+1}X_{t}X_{t}^{\top}\right){\theta^{\star}}+\frac{d}{2}\log\left(1+\frac{1}{8\alpha}\right)$
@@
$\mathbb{E}_{S_{T},S_{T}^{\prime},(X,Y)}[-\log\widetilde{p}(X,Y)]$ $\leq\frac{1}{T+1}\mathbb{E}\left[-\sum\limits_{t=1}^{T+1}\log\tilde{\rho}_{t}(X_{t},Y_{t})\right]$
$\leq\mathbb{E}_{(X,Y)}[-\log\sigma(Y\langle X,{\theta^{\star}}\rangle)]+\alpha\lambda_{\max}\left(\mathbb{E}[XX^{\top}]\right)\|{\theta^{\star}}\|^{2}+\frac{4d\log(2T)}{T+1}$
$\leq\mathbb{E}_{(X,Y)}[-\log\sigma(Y\langle X,{\theta^{\star}}\rangle)]+\frac{5d\log(2T)}{T},$
where the last inequality uses 29. For the second case, suppose $\|{\theta^{\star}}\|>T^{4}$. For this, we have
@@
[…]
$\quad\quad\quad\quad\quad+\mathbbm{1}\left(\forall t,Y_{t}=\operatorname{sign}(\langle X_{t},{\theta^{\star}}\rangle),\mathbb{P}_{X,Y\sim\mathcal{D}_{{\theta^{\star}}}}\left(\langle X,{\theta^{\star}}\rangle\langle X,\hat{\theta}\rangle<0\right)>\frac{C_{1}(d+\log T)}{T}\right)$
$\left.\left.\quad\quad\quad\quad\quad+\mathbbm{1}\left(\mathbb{P}_{X,Y\sim\mathcal{D}_{{\theta^{\star}}}}\left(\langle X,{\theta^{\star}}\rangle\langle X,\hat{\theta}\rangle<0\right)\leq\frac{C_{1}(d+\log T)}{T}\right)\right)\right]$
@@
Then, by Lemmas 33, 31 and 32, we have
@@ 7 Directions of Future Research
  • Our online-to-batch conversion argument in Section 6 only provides in-expectation results in important applications, such as logistic regression. While recent martingale-based arguments allow for high-probability excess risk bounds (see, e.g., (van der Hoeven, Zhivotovskiy, and Cesa-Bianchi, 2023)), these existing arguments do not extend straightforwardly to the transductive setup, where the natural martingale structure is not present. An even more challenging question would be to provide computationally efficient methods for logistic regression with no assumptions on the design, for which the excess risk bound of Corollary 26 holds with high probability.
@@
  • Our results currently apply only to linear classes; a natural next step is to extend them to more general classes. At present, however, we lack the technical tools to construct natural, nontrivial priors for such classes, beyond, for example, the uniform distribution on $\varepsilon$-covers.
  • While some of our techniques, such as exponential weights and mixability, extend to broader loss families, our analysis also exploits specific structural properties of the considered models, including the form of the sigmoid link and the resulting Hessian in the parameter space. For instance, our logistic regression results rely on the slab construction and need not hold for the same logarithmic loss in more general GLM settings. As in the case of general classes, it would be interesting to study other losses and different link functions in the context of our setup.
Acknowledgments and Disclosure of Funding
JQ and AR acknowledge support from ARO through award W911NF-21-1-0328, as well as the Simons Foundation and the NSF through awards DMS-2031883 and PHY-2019786. The authors thank Jaouad Mourtada for identifying an issue with the polynomial-time implementation of exponential weights with heavy-tailed priors in the initial version of the paper and for contributing to the discussion on the related sampling strategy described in (Dalalyan and Tsybakov, 2012b). The authors also thank Jingfeng Wu for pointing out the reference (Hsu and Mazumdar, 2024). Finally, the authors are thankful to the referees whose precise comments improved the presentation of the paper.
@@ Appendix A Additional Results for Sparse Regression
Before discussing these discrete priors, we note that despite the minor penalty of the logarithm of the inverse smallest scaled singular value constant $\kappa_{s}$, there is a significant advantage: the predictor in Theorem 3 has an explicit Gaussian-type integral form. On the other hand, the predictor derived from sparsity priors can be quite challenging to exploit, as it might require too many samples from this prior (or equivalently, sampling from the posterior distribution $\rho_{t}$ is hard). However, from a practical implementation standpoint, it can still be effective, as claimed by simulations in (Rigollet and Tsybakov, 2011, 2012).
@@
where $H_{d}=\sum\nolimits_{i=1}^{d}\exp(-i)$. Although this prior is not data-dependent, it can be naturally combined with the data-dependent prior (6) by introducing a product prior that first samples the sparsity pattern vector according to (45), then uses (6) restricted to the coordinates corresponding to this sparsity pattern. Finally, the exponential weights algorithm is run with respect to this data-dependent prior. To slightly simplify the analysis, we will use a version of Theorem 1, where the additional clipping step is performed, and then we use aggregation using the exp-concavity of the squared loss in the bounded case.
@@
$\widetilde{\theta}_{t,p}=\arg\min_{\theta\in\Theta_{p}}\left\{\sum_{i=1}^{t-1}(y_{i}-\langle x_{i},\theta\rangle)^{2}+\lambda\sum\limits_{t=1}^{T}(\langle x_{t},\theta\rangle)^{2}\right\},$
@@
The work of Rigollet and Tsybakov (2012) suggests a Metropolis–Hastings type algorithm for such sampling, though in the worst case its convergence rate can be quite slow.
@@
Proof. Fix any $p\in\mathcal{P}$ such that $|p|=s$. When restricted to the coordinates corresponding to the set $p$, the bound of Corollary 41 in Appendix C gives the regret bound
@@
It remains to notice that for any $p\in\mathcal{P}$ such that $|p|=s$, the prediction $\operatorname{clip_{m}}\left(\langle x_{t},\widetilde{\theta}_{t,p}\rangle\right)$ is absolutely bounded by $m$. Finally, combining the exp-concavity of the squared loss in the bounded case (Vovk, 2001, Remark 3) (thus, we choose $\eta=\frac{1}{8m^{2}}$ in the exponential weights algorithm) and the fact that the predictor (46) is a posterior average with respect to the exponential weights algorithm with prior $\mu$ over $\mathcal{P}$, and by Lemma 5 with $\log\left(\frac{1}{\mu(p)}\right)=2s\log\left(\frac{ed}{s}\right)+\frac{1}{2}$ by (Rigollet and Tsybakov, 2012, Inequality (5.4)), we obtain that for any $s$-sparse $\theta^{\star}\in\mathbb{R}^{d}$,
@@ Appendix B Mixability of the Squared Loss for Unbounded Sets of Predictors
Lemma 35 (Restatement of Lemma 6).
@@
satisfies $\frac{\partial^{2}h(y)}{\partial y^{2}}\geq 0$, making $h$ convex. Importantly, the derivation of $\frac{\partial^{2}h(y)}{\partial y^{2}}$ does not impose any boundedness condition on $f_{\theta}(x)$. Since $h$ is convex, its maximum is attained at either 1-1 or $1$. Therefore, by symmetry, $\widehat{f}$ satisfies
@@ Appendix C Revisiting Existing Bounds for Sequential Linear Regression
Our loss function is the logarithmic function, and we simply compute the total error for the mix-loss as suggested by Lemma 4. This straightforward computation is sufficient to recover all existing sequential online bounds, including the one in (Vovk, 2001), the bound for clipped ridge regression in (Kivinen and Warmuth, 1999), and the bound of Gaillard et al. (2019), which has already been discussed in Section 2. In the following, we discuss how essentially the same strategy, specific to sequential ridge regression, was previously employed by Zhdanov and Vovk (2009). Here, we mainly provide a detailed exposition of this approach and extend it to our transductive setup.
@@
Proposition 36 (Generalization of (Azoury and Warmuth, 2001; Zhdanov and Vovk, 2009)).
@@
Although the bound of Proposition 36 can be derived directly using linear algebraic arguments (which is essentially done in the proof of Theorem 4.6 in (Azoury and Warmuth, 2001)), we believe that the following proof strategy is more insightful. It provides this identity without prior knowledge of its possible existence. We note that the proof strategy and explicit identity in Proposition 36 closely align with previous work, specifically with Zhdanov and Vovk (2009, Theorem 1), which establishes this identity for $A=I_{d}$ and acknowledges that a similar identity appears in Azoury and Warmuth (2001). Additionally, the work by Zhdanov and Kalnishkan (2013) explores sequential formulations of kernel ridge regression with the focus on similar identities. Although this formulation is not novel, it is included here for completeness and to support the transductive prior framework used in this paper. The extension to a general positive definite matrix $A$ as the regularizer corresponds directly to the transductive priors considered, simplifying derivations of bounds such as that in Gaillard, Gerchinovitz, Huard, and Stoltz (2019), as well as additional technical results used throughout the paper.
@@
$\log\left(\mathbb{E}\exp(-(y-Z)^{2}))\right)=\frac{(\nu-y)^{2}}{2\sigma^{2}+1}+\frac{1}{2}\log(2\sigma^{2}+1).$
Proof. Using Lemma 16, we have
$-\log\left(\mathbb{E}\exp(-(y-Z)^{2}))\right)$ $=-\log\left(\frac{1}{\sqrt{2\pi}\sigma}\int\exp(-(y-z)^{2}-(z-\nu)^{2}/2\sigma^{2})dz\right)$
@@
$=-\log\left(\frac{1}{\sqrt{2\sigma^{2}+1}}\exp\left(-\frac{(\nu-y)^{2}}{2\sigma^{2}+1}\right)\right)$
@@
We are ready to provide a proof of Proposition 36.
@@
We want to compute the right-hand side of (11) for the class (48). The first computation uses Lemma 16. We have
@@
Another perspective on (49) is the following. We assume that at step $t$ we predict with the distribution $\rho_{t}$ over $\mathbb{R}^{d}$ using the exponential weights over the class (48) with respect to the logarithmic loss. By Lemma 16, we have
$\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\limits_{t=1}^{T}\pi(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right)=-\sum\limits_{t=1}^{T}\log\left(\mathbb{E}_{\theta\sim\rho_{t}}\exp\left(-\pi(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right).$ (51)
@@
Clearly, the distribution $\rho_{t}$ is multivariate Gaussian. Using Lemma 16, we can immediately obtain
@@
Observe that if $\theta\sim\rho_{t}$, then $\langle x_{t},\theta\rangle\sim\mathcal{N}\left(\langle\widehat{\theta}_{t},x_{t}\rangle,\frac{1}{2\pi}x_{t}^{\top}\left(\sum\limits_{i=1}^{t-1}x_{i}x_{i}^{\top}+A\right)^{-1}x_{t}\right)$, which implies by Lemma 37 that the following holds
@@
Finally, we have by Lemma 38
@@
$-\log\left(\mathbb{E}_{\theta\sim\mu}\exp\left(-\sum\limits_{t=1}^{T}\pi(y_{t}-\langle x_{t},\theta\rangle)^{2}\right)\right)$
$=\sum\limits_{t=1}^{T}\left(\pi(1-h_{x_{t}})(\langle\widehat{\theta}_{i},x_{t}\rangle-y_{t})^{2}+\frac{1}{2}\log\left(\frac{1}{1-h_{x_{t}}}\right)\right).$ (52)
Denote $A_{t}=x_{t}^{\top}\left(\sum\limits_{i=1}^{t-1}x_{i}x_{i}^{\top}+A\right)^{-1}x_{t}$. Finally, by Lemma 38 we have
$\sum\limits_{t=1}^{T}\frac{1}{2}\log\left(\frac{1}{1-h_{x_{t}}}\right)=\frac{1}{2}\sum\limits_{t=1}^{T}\log\left(\frac{\det(A_{t+1})}{\det(A_{t})}\right)$
$=-\frac{1}{2}\log(\det(A))+\frac{1}{2}\log\left(\det\left(\sum\limits_{t=1}^{T}x_{t}x_{t}^{\top}+A\right)\right).$ (53)
where in the last computation, we used the telescopic sum. Combining (50), (C), and (C), we prove the claim. ∎
The result of Proposition 36 immediately implies the following classical result.
Corollary 39 (The regret of the non-linear predictor of Vovk (2001)).
@@
Proof. To see how the implication, one should first verify, using Lemma 37, that
@@
Combining this with Proposition 36, we have